### v2

# MDA: An Interpretable Multi-Modal Fusion with Missing Modalities and Intrinsic Noise

Abstract
Multi-modal fusion is crucial in medical data research, enabling a comprehensive understanding of diseases and improving diagnostic performance by combining diverse modalities. However, multi-modal fusion faces challenges, including capturing interactions between modalities, addressing missing modalities, handling erroneous modal information, and ensuring interpretability. Many existing researchers tend to design different solutions for these problems, often overlooking the commonalities among them. This paper introduces a novel multi-modal fusion framework that introduces Modal-Domain Attention (MDA), which constructs linear attention relationships between modalities through continuous attention, thereby achieving adaptive adjustment of the weights for each modality. It aims to facilitate the fusion of multi-modal information while allowing for the inclusion of missing modalities or intrinsic noise, thereby enhancing the representation of multi-modal data. We provide visualizations of accuracy changes and MDA weights by observing the modal fusion process, offering a comprehensive analysis of its interpretability. Experimental results indicate that the proposed MDA approach maintains state-of-the-art (SOTA) performance across various tasks on multiple public datasets. Worth noting is that we will make publicly available a multi-modal database for gastrointestinal diseases, which includes WLE, EUS, and report data. The visualization outcomes of MDA within the gastrointestinal disease dataset align closely with the conclusions of existing clinical studies regarding the dependency of different diseases on multiple modalities. The associated code will also be released to facilitate future research endeavors.
Introduction
Recently, multi-modal fusion learning has demonstrated outstanding performance across various tasks (Huang et al. 2021), particularly in medical applications, as it integrates medical information from different modalities, offering comprehensive diagnostic evidence for healthcare professionals (Litjens et al. 2017). It has been employed in tasks such as disease-assisted diagnosis (Manh et al. 2022; Jiao et al. 2024) and prognostic forecasting (Ren et al. 2024a).
Despite efforts in medical multi-modal research, fundamental and challenging issues remain due to the complexity of multi-modal clinical data and the real-world application scenarios: Chanllenge1: The difficulty of multi-modal fusion arising from modality heterogeneity. Despite the individual strengths of each modality, integrating different modalities, including images and text, to improve disease diagnosis is still hindered by modality heterogeneity, which arises from the diverse data representations across different modalities. The cross-attention method is considered an efficient multi-modal fusion technique due to its ability to better capture the complex interrelationships between modalities, which have been applied to tasks such as image recognition diagnosis (Sun et al. 2023; Ma et al. 2023), object detection (Shen et al. 2024; Wen et al. 2024), segmentation (Ates, Mohan, and Celik 2023; Liu et al. 2024; Lin et al. 2023), and image fusion (Li and Wu 2024), among others. However, due to the adherence to the mutual attention between dual modalities, they are challenging to extend to tri-modal or even more modalities. Chanllenge2: The issue of modality missing in real-world scenarios. Missing modalities have garnered widespread interest in the medical fields (Wang et al. 2023; Ma et al. 2021). Methods for handling missing modalities can be broadly categorized into two types (Zhang et al. 2024): inference-based and non-inference-based. Inference-based methods (Dong and Sun 2023; Yang et al. 2022; Li et al. 2023; Ren et al. 2024b) strive to generate information for the missing modalities, however, they often overlook the assessment of the quality of the generated missing modalities and may lead to model collapse (Shumailov et al. 2024). For non-inference-based approaches (Liu et al. 2023; Xu et al. 2022b; Zhang et al. 2019; Wen et al. 2023) to incomplete multi-modal learning, they focus more on utilizing the available modalities to explore information and understand multi-modal data, but the information gap caused by missing modalities is difficult to bridge. Chanllenge3:Intrinsic noise. Multimodal learning that incorporates text may encounter the issue of intrinsic noise, which is more prevalent in the medical domain due to the heavy reliance on subjective annotations by clinical experts in diagnostic reports (Fusaroli et al. 2021). Intrinsic noise refers to the hidden inherent noise present in the text, which contradicts the knowledge conveyed by other modalities and may influence the final predictions. For example, it could occur when the report describes the echogenicity of a tumor in an EUS image as "hyperechoic" while the image itself reveals that the echogenicity is "heterogeneous." Selectively filtering accurate and advantageous multimodal information while avoiding errors that could compromise diagnostic accuracy remains an unresolved research challenge in multimodal feature fusion. Chanllenge4: The comprehensive interpretability of multi-modal fusion. Correctly interpreting AI model decisions has been a persistent challenge in achieving interpretability in multi-modal fusion (Huff, Weisman, and Jeraj 2021; Salahuddin et al. 2022). Currently, the interpretability of medical-assisted diagnostic models primarily focuses on generating heatmaps (Selvaraju et al. 2017; Wang et al. 2020) to visualize the network’s attention and enhance model interpretability (Manh et al. 2022; Wu et al. 2023). However, heatmaps are visualizations of network parameters and cannot explain modality specificity in different cases, i.e., they cannot reflect modality weights for different cases. In the context of diagnosing gastrointestinal diseases, different cases may focus on different modalities (Jacobson et al. 2023), and the inability to capture the inconsistency between modalities or decisions that contradict clinical experience may lack credibility. The interpretability of fine-grained features is undoubtedly crucial. However, the interpretability of modalities, that is, discussing the gold standard modalities for diagnosing various diseases, is equally important  (Richardson et al. 1995; Chandler et al. 2019; Jacobson et al. 2023). To address the aforementioned challenges, we propose a novel modality fusion approach wherein we introduce MDA. The objective of MDA is to compute the saliency of a particular modality while considering the interactive influences of all other modalities, thereby achieving a comprehensive understanding of the multimodal information within the overall scene. This is primarily achieved through the implementation of a novel continuous attention algorithm, which constructs linear attention relationships between multi-modal features, thereby dynamically adjusting the attention allocation between different modalities for the thorough integration of multi-source heterogeneous information (as shown in Fig. 1). Furthermore, this approach ensures the scalability of the fusion framework, offering a new solution for scalable multimodal fusion. Additionally, it pioneers the exploration of the macroscopic interpretability of multimodal learning.
Figure 1: A unified multi-modal learning strategy involves learning with different multi-modal configurations. (a) Train and test with full modality. (b) The model will reduce its attention when learning with missing modalities or intrinsic noise.
  • We propose a novel multi-modal fusion framework that integrates MDA, which constructs linear relationships between multi-modal information through continuous attention, thereby achieving adaptive control over the weights of multiple modalities and demonstrating modality scalability.
  • MDA provides a unified solution to several challenges in multi-modal fusion, including multi-modal integration, modality missing, and learning with intrinsic noise.
  • Supplementary experiments indicate that we achieve SOTA performance across multiple tasks on various public datasets. Additionally, we provide a publicly accessible multi-modal gastrointestinal stromal tumor dataset.
  • Based on MDA, we comprehensively analyze the sources of multi-modal efficacy from a macroscopic perspective (across different diseases) and a microscopic perspective (within individual cases) while providing interpretability to the model from a clinical experience perspective.
Figure 2: Overview of the proposed network framework. The uni-modal training involves building pre-trained models for multi-disease classification across different modalities, with a dedicated classifier assigned to each disease. The work introduces the weight calculation for each scenario under the proposed multi-modal attention module.
Methodology
This study aims to enhance collaborative knowledge learning capabilities in various modalities, including different image modalities, text modality, audio modality, and others. Considering there are ${M}$ image modalities ${I}$, ${P}$ text modalities ${T}$, and ${Q}$ audio modalities ${A}$, the total number of modalities is ${N(N=M+P+Q)}$. The multi-modal data ${\mathcal{X}}$ can be represented as $\mathcal{X}=\{\{\mathbf{I}_{j}\}_{j=1}^{M},\{\mathbf{T}_{k}\}_{k=1}^{P},\{\mathbf{A}_{l}\}_{l=1}^{Q}\}$ and the predicted labels as ${Y}$.
An overview of the proposed method is shown in Fig. 2. we first construct individual classification tasks ${\mathcal{T}_{j}=\{\mathbf{I}_{j},Y\}}$ for modal ${j}$, the single-task definitions for other modalities are similar. Then, we feed the extracted features into the MDA to calculate the adaptive weights for each modality. Finally, we perform multi-modal feature fusion based on the inter-modality weights and feed the fused features into the classifier to accomplish the classification task. The entire multi-modal classification task be defined as ${\mathcal{T}=\{\mathcal{X},Y\}}$, ${Y}$ is represented as a one-hot vector for this multi-class task. Next, we will define the processing approaches for different modalities.
Image modalities
Pre-trained convolutional neural networks (CNNs) is utilized as feature extractors for image modalities, such as ResNet. Take image modality $\mathbf{I}_{j}$ as an example, the vectorized latent features ${\mathbf{f}_{j}}$ obtained from the feature extractors ${\mathbf{G}_{j}}$ are denoted as:
${\mathbf{f}_{j}}=\mathbf{G}_{j}\left(\mathbf{I}_{j}\right),$ (1)
Multi-modal data fusion requires attention to inter-modal learning mechanisms and individual modality learning efficiency, both crucial for effective data fusion. We introduce self-attention to enable deeper feature exploration, enhancing cross-modal fusion and complementing inter-modal learning. For image features, the calculation of self-attention $\mathbf{SA}$ scores is as Eq. 2, where ${Q_{j}}$, ${K_{j}}$, and ${V_{j}}$ are linear transformations of the latent feature $\mathbf{f}_{j}$, ${\mathbf{\Phi}_{j}}$ represents the self-attention features of the ${j}$ modal.
$\mathbf{\Phi}_{j}=\mathbf{SA}(\mathbf{f}_{j})=\operatorname{softmax}\left(\frac{Q_{j}\cdot(K_{j})^{T}}{\sqrt{d}}\right)\cdot V_{j}$ (2)
Text modality
We employ BERT (Devlin et al. 2018) to encode the input text. BERT learns contextual representations of words or subwords in a text by using a self-attention mechanism. This is consistent with the self-attention module in image feature extraction. To adapt BERT to our specific task, we unfroze the last four layers of BERT for training. The self-attention feature extraction $\mathbf{\Phi}_{k}$ representation for textual reports $\mathbf{T}_{k}$ is as follows:
$\mathbf{\Phi}_{k}=\mathbf{BERT}\left(\mathbf{T}_{k}\right)$ (3)
Audio modality
We utilize Mel-frequency cepstral coefficients (MFCCs) (Tzanetakis and Cook 2002) as the representation for the audio modality. MFCCs have gained extensive application in the field of audio and speech processing due to their ability to mimic human auditory perception, facilitate dimensionality reduction of features, exhibit robustness to variations, and demonstrate broad applicability. We also applied self-attention to the extracted audio features, take audio modality ${\mathbf{A}_{l}}$ as an example, and its feature representation $\mathbf{\Phi}_{l}$ obtained after processing through the self-attention mechanism is denoted as:
$\mathbf{\Phi}_{l}=\mathbf{SA}(\mathbf{f}_{l})=\mathbf{SA}(\mathbf{MFCCs}\left(\mathbf{A}_{l}\right))$ (4)
Continuous attention mechanism
Existing multi-modal attention mechanisms are typically designed for bi-modal fusion, focusing on point-to-point interactions between modalities. However, extending these methods to handle three or more modalities leads to computational complexity and a lack of consideration for interactions between modalities. In contrast, MDA aims to compute the attention of each modality while considering the influences of all other modalities, enabling a comprehensive understanding of multi-modal information. It achieves this by employing continuous attention mechanisms to dynamically adjust the attention across all modalities and to establish linear relationships between attentions, facilitating rapid and dynamic adjustments. The continuous attention mechanism employed in MDA constructs linear attention relationships among multiple modalities. The specific computational approach is illustrated as Eq. 5, where $\mathbf{\Theta}_{i}^{N-1}$ represents the attention value computed between modality $i$ and the other $N-1$ modalities, $R$ denotes the ReLU activation.
$\mathbf{\Theta}^{e}_{i}=\begin{cases}\operatorname{R}\left(\frac{Q_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},e=1\\ \operatorname{R}\left(\frac{\mathbf{\Theta}^{e-1}_{i}\cdot(K_{e})^{T}}{\sqrt{d}}\right)\cdot V_{i}^{e},1\leq e\leq N-1\end{cases}$ (5)
MDA establishes a linear relationship between multi-modal attentions through continuous attention calculations, allowing for infinite expansion without being limited to handling only two or three modalities. This scalability is crucial for applications that require continuous integration of new modalities over time, particularly in medical imaging, where the collection of more modal data is necessary to provide accurate and trustworthy judgments.
Objective function
For the classification training of single-modality model $i$, given the features ${\mathbf{\Phi}_{i}}$ obtained after the self-attention module, we employ a multi-layer perceptron classifier ${\mathcal{C}_{s}}$ for disease prediction. The model training is guided by the cross-entropy loss, defined as Eq. 6. For the multi-modal classification training, the classification loss is defined as Eq. 7, where ${\mathcal{C}_{f}}$ stands for multi-modal classifier.
$\ell_{i}=\text{ CrossEntropy }\left(\mathcal{C}_{s}\left(\mathbf{\Phi}_{i}\right),Y\right)$ (6)
= CrossEntropy ($\ell=\text{ CrossEntropy }\left(\mathcal{C}_{f}\left(\mathbf{\{}\Theta_{i}^{N-1}\}_{i=1}^{N}\right),Y\right)$ (7)
Training setting
We have directly developed a model that operates under conditions of missing modalities and intrinsic noise. We processed the individual modality features obtained to simulate these scenarios. Specifically, within each batch of data, we randomly selected ${(N+P)}$ samples and subjected them to the following treatments: $N$ modality’s features were sequentially set to 1 to simulate missing modalities, and one random attribute in $P$ textual modality was replaced with another random label to simulate intrinsic noise.
We employ the Adam optimizer with a weight decay rate of 1e-4 to optimize the model parameters. All experiments were conducted using the PyTorch framework with the Geforce RTX 4090 GPU.
Datasets
We aim to develop a general multi-modal fusion model that can effectively handle modality heterogeneity, adapt to varying degrees of missing modalities, filter out inconsistent and intrinsic noise, and mitigate their negative impact on the results. To this end, we conducted experiments using Four different multi-modal datasets, encompassing diverse dimensions such as medical data, natural images, movies, and audio, to assess the generality of the proposed model comprehensively.
Gastrointestinal disease dataset We compiled a multi-center gastrointestinal disease dataset comprising three modalities: Endoscopic Ultrasound (EUS), White Light Endoscopy (WLE), and imaging reports. The dataset was gathered from two separate centers (***hospital and ***hospital) and captured using Olympus UM-2R and UM-3R equipment, totaling 421 cases spanning eight distinct categories (Gastrointestinal Stromal Tumor (GIST), Gastric Leiomyoma (GLM), Neuroendocrine Tumor (NET), Ectopic Pancrea (EP), Lipoma, Gastrointestinal Schwannomas (GS), Pneumatosis Cystoid (PC), and Inflammatory Fibroid Polyp (IFP)). Among these patients, the gender distribution is 53% male to 47% female, with an age range spanning from 27 to 78 years. The imaging reports offer detailed descriptions of lesion characteristics observed within the two imaging modalities. The detailed data quantities for each category are shown in Table 1.
Table 1: Detail of the Gastrointestinal disease dataset.
# Cases # EUS # WLE # Reports
GIST 149 1539 460 149
GLM 117 1123 340 117
NET 107 1187 404 107
EP 18 182 65 18
Lipoma 16 180 61 16
GS 5 77 18 5
PC 6 65 14 6
IFP 3 24 8 3
Overall 421 4213 1370 421
Derm7pt dataset (Kawahara et al. 2018) is a multi-modal dataset that contains clinical, dermoscopic, and meta-data modals for melanoma skin lesions in dermatology. We predict disease diagnosis (DIAG) and all 7 dermoscopic concepts, including blue-whitish veil (BWV), dots and globules (DaG), pigmentation (PIG), pigment network (PN), regression structures (RS), streaks (STR), and vascular structures (VS) in considered datasets. The dataset contains a total of 1011 cases.
Multimodal IMDb (MM-IMDb) (Arevalo et al. 2017) contains two modalities: image and text. We conduct experiments on this dataset to predict a movie genre using image and text modality, which is a multi-label classification task as multiple genres could be assigned to a single movie. The dataset includes 25, 956 movies and 23 classes.
Audiovision-MNIST (Vielzeuf et al. 2018) consists of an independent image and audio modalities. The images, which are digits from 0 to 9, are collected from the MNIST dataset (LeCun et al. 1998) with a size of 28 × 28, and the audio modality is collected from Free Spoken Digits Dataset containing raw 1500 audios. The dataset contains 1500 samples for both image and audio modalities.
Table 2: Multi-modal Fusion Results Under Perturbed Inputs with Ablation Experiments.
Type Modal Fusion Method Accuracy
Pre-trained Uni-modal EUS - 85.3
Uni-modal WLE - 81.7
Uni-modal Report - 87.4
Full Modalities Multi-modal EUS+WLE+Report w/o MDA 91.2
Multi-modal EUS+WLE+Report SCT Fusion (2023) 82.8
Multi-modal EUS+WLE+Report w/ MDA (Ours) 98.9
Missing Modalities Multi-modal Randomly missing 1 modal w/o MDA 88.1
Multi-modal Fixed missing 1 modal MMD (2022) 88.4
Multi-modal Randomly missing 1 modal w/ MDA (Ours) 95.4
Learning with Intrinsic Noise Uni-modal Report - 80.5
Multi-modal EUS+WLE+Report w/o MDA 89.3
Multi-modal EUS+WLE+Report w/ MDA (Ours) 97.9
    We simulated missing modalities by applying dropout to three different modalities. The full set of modalities was used in training, while during testing, only the modality with dropout was missing. The final result was obtained by averaging the predictions from the three models.
Experiments and discussion
In the first section, we present the results of our proposed method in addressing the challenges of modality heterogeneity, missing modalities, and learning with intrinsic noise in our gastrointestinal disease dataset. We comprehensively validate the effectiveness of our proposed method by analyzing three different fusion scenarios and the results of the ablation experiment. In the second section, We conducted experiments on publicly available datasets to evaluate the performance of the proposed method in multi-modal fusion and fusion efficacy in the presence of missing modalities. We compared our method with SOTA approaches for handling missing modalities, and in multiple tasks across multiple databases, MDA achieved SOTA results. The interpretability analysis of MDA is presented in the last section.
The efficacy of MDA in confronting the three key challenges of multi-modal fusion
Classifying Gastrointestinal tumors is crucial for surgical decision-making in clinical (Liu et al. 2021). Given their multi-modal characteristics, this paper will use submucosal tumor classification as an example to demonstrate our methodology. We evaluated the impact of EUS, WLE, and imaging reports in multi-modal fusion under three challenges, examining performance variations and the importance of each component in conducting ablation studies (see Table 2).
Modality Heterogeneity Challenge. The training of single-modal tasks is independent, with uni-modal features integrated into the MDA after model freezing. This is done to ensure that the self-attention module within the single-modal model does not engage in the learning process of the feature fusion module MDA, thereby eliminating any ambiguity in the analysis of MDA’s performance enhancement, as it does not affect the weight adjustments of MDA. The experimental results show that the direct concatenation-based fusion of multi-modal features achieved higher accuracy than any individual uni-modal testing accuracy, this indicates the effectiveness of multi-modal fusion in improving disease diagnosis accuracy on this dataset. By employing the MDA, which adaptively learns the weight relationships between modalities, the fusion capability of the multi-modal model was enhanced, resulting in a significant improvement in diagnostic accuracy compared to direct concatenation (acc: 91.2% $<$ 98.9%). This demonstrates the effectiveness of the proposed model in addressing modality heterogeneity.
The precision of MDA significantly surpasses that of other scalable multimodal methods such as SCT, demonstrating its advantage in handling mutual attention across modalities. MDA employs continuous attention learning without the need for additional branches, whereas SCT incurs a heavy computational burden and lacks sustained attention in fusion, resulting in suboptimal integration performance. Sustained attention is also a crucial factor that enables MDA to surpass the sustainable scalability in multimodality without augmenting the computational load of branching pathways.
Missing modality challenge. The results showed that the accuracy of direct concatenation fusion was lower than the accuracy of concatenation fusion using all modalities (acc: 88.1% $<$ 91.2%), indicating that the missing modality indeed affects the performance of multi-modal fusion, even with partial missing. However, when using MDA, even with only two modalities involved, the accuracy under attentional adjustment far exceeds that obtained with concat (accuracy: 95.4% $>$ 88.1%). This finding suggests that MDA can counteract most of the negative effects of missing modalities on feature fusion by shifting attention, as evidenced by the fact that it is only about 3% lower than the fusion accuracy using all modalities (acc: 95.4% $<$ 98.9%).
Learning with intrinsic noise challenge. The results showed that the wrong attribute significantly impacted the diagnostic accuracy of the model, as the accuracy decreased from 87.4% to 80.5%. After applying the concatenation-based multi-modal fusion, the accuracy improved to 89.3%. Although the image information helped mitigate the impact of erroneous modalities, the accuracy was still lower than that achieved with correct modalities (acc: 89.3% $<$ 91.2%). Upon incorporating the MDA, the model is capable of disregarding misleading information, as it establishes a correspondence between the report and image modalities. As a result, the diagnostic accuracy was essentially restored to the level achieved with the correct modalities (acc: 97.9%).
We compared the recent multi-modal fusion methods on a gastrointestinal dataset under the full modality setting. The results are shown in Table 2. For the SCT Fusion method (Hoffmann, Clasen, and Demir 2023), we computed the cross-entropy loss between our classification results and the ground truth of the eight-class labels without modifying the other configurations. The results demonstrate that the proposed method significantly outperforms recent multi-modal approaches in terms of fusion performance across multiple modalities (acc: 98.9% $>$ 82.8%). We observe that a pivotal reason for this outcome is the specialized nature of these methods, which have been designed for particular tasks, leading to reduced model portability. In contrast, the design flexibility of MDA enables its application across a broad spectrum of multimodal datasets.
Table 3: Comparison of classification results to the state-of-the-art multi-modal fusion methods in terms of accuracy (%) by using Derm7pt (Kawahara et al. 2018) dataset.
Method BWV DaG PIG PN RS STR VS DIAG avg.
Baseline (Kawahara et al. 2018) 87.1 60.0 66.1 70.9 77.2 74.2 79.7 74.2 73.7
HcCNN (Bi et al. 2020) 87.1 65.6 68.6 70.6 80.8 71.6 84.8 69.9 74.9
FM4-SS (Tang et al. 2022) 88.1 66.1 70.1 71.1 81.5 78.0 81.8 78.5 77.0
RemixFormer (Xu et al. 2022a) - - - - - - - 81.3 -
CAFNet (He et al. 2023) 87.8 61.5 73.4 70.1 81.8 77.0 84.3 78.2 76.8
TFormer (Zhang, Xie, and Chen 2023) 86.7 66.9 70.3 74.3 82.1 76.7 83.0 79.5 77.4
MICA (Bie, Luo, and Chen 2024) 84.4 79.1 68.8 74.4 72.8 71.3 81.3 83.9 77.0
RGMC50 (Li et al. 2024) - - - - - - - 78.9 -
MDA (Ours) 91.6 66.9 70.4 72.4 79.8 75.8 85.2 85.7 78.5
Table 4: Model performance comparison of classification accuracy of missing modality (by setting different available audio rates) on avMNIST dataset. And train 30 epochs. ’M-m Concat’ refers to multi-modal fusion using concatenation, while ’Ours-m’ represents our proposed fusion method that handles missing modalities during testing. ’Ours-f’ denotes testing with full modalities, while only having nof the audio modality available in the training data.
Multi-Modal Rate Uni-Image Uni-Audio M-m Concat Ours-m SMIL (2021) ShaSpec (2023) Ours-f
70% 87.3 82.5 87.9 95.8 96.4 97.3 98.1
50% 87.3 82.5 84.3 86.3 95.6 96.0 96.2
20% 87.3 82.5 83.6 86.7 94.4 95.3 95.9
    Results for this case where only audio data is missing. Our approach involves random modalities being missing.
Comprehensive comparison in different database
In this chapter, we primarily utilize open datasets to conduct a comprehensive comparison of the proposed MDA method with the latest multi-modal fusion techniques across three major challenges. In the first section, we compared our MDA method with the latest techniques on the Derm7pt dataset, using a full-modal approach. In the second section, we compared MDA with the latest methods on two public multimodal datasets under the scenario of missing modalities.
Comparison with related multi-modal fusion works To comprehensively evaluate MDA, we compared it with other state-of-the-art multi-modal fusion methods on the Derm7pt dataset, as shown in Table 3. The table reports the quantitative results of the classification accuracy for each concept, highlighting the performance of MDA on various tasks. MDA outperforms other methods in most metrics, particularly in disease diagnosis, where it achieves an accuracy of 85.7%, significantly surpassing the second-best method by 1.8%. This exceptional performance in disease diagnosis demonstrates the robustness and effectiveness of MDA in integrating and analyzing multi-modal data, which is crucial for accurate diagnosis in medical studies.
Table 5: Model performance comparison of classification accuracy of missing modality (by setting different available audio rates) on MM-IMDb dataset. Evaluating performance using F1 sample scores and training 5 epochs.
Multi-Modal Rate Uni-Image Uni-Text M-m Concat SMIL Ours-m
70% 0.361 0.398 0.637 - 0.701
50% 0.361 0.398 0.533 - 0.578
20% 0.361 0.398 0.496 0.541 0.546
    Results for this case where only text data is missing. Our approach involves random modalities being missing.
Comparison with related missing modality works Due to the extensive research on missing modalities, we specifically conducted an extended investigation into the performance of MDA in the presence of missing modalities. Experiments were performed on the MM-IMDb and avMNIST datasets. We constructed datasets with 20%, 50%, and 70% of the full modality and randomized the missing modalities during training. We compared the accuracy of single-modal, direct multi-modal fusion, and the proposed method during evaluation and testing. We also tested the proposed model’s proficiency in multi-modal fusion using 100% images and naudio during training, comparing the results with SMIL (Ma et al. 2021) and ShaSpec (Wang et al. 2023).
Table 4 and Table 5 present experimental results on avMNIST and MM-IMDb datasets. The results indicate that when only 20% of the multimodal data is included, MDA may underperform compared to single-modality results due to the lack of sufficient information to learn the linear attention relationships between the two modalities. However, the results are still superior to concatenation-based fusion. When evaluating the influence of varying multi-modal rates on accuracy, MDA exhibits a markedly superior resistance to interference compared to the concatenation fusion method. This suggests that MDA is proficient in leveraging limited information, which is a design strength of MDA, enabling the adaptive allocation of weights to information from different modalities. Table 4 presents the performance of Ours-f when tested with both audio and image modalities, with only nof the audio available in training. MDA outperforms SMIL and ShaSpec, especially when 70% of the audio is missing. This is because MDA’s ability to adaptively transfer attention from missing modalities to others preserves feature reliability, whereas other methods may compromise feature quality through generated missing modalities. Moreover, MDA’s versatility in handling any missing modality makes it more applicable to real-world scenarios.
Figure 3: Macroscopic investigation of attention weights for various diseases
Interpretability analysis
In medical research, attention visualization methods like CAM are used for interpretability but have limitations in analyzing model attention across modalities and disease categories. To address this, we use adaptive attention maps across modalities to comprehensively analyze model performance in multi-modal fusion, modality absence, and handling intrinsic noise. Interpretability in Diagnostic Standards Research In clinical practice, multi-modal imaging techniques are frequently employed to gather comprehensive diagnostic information. For instance, the combination of EUS, WLE, and CT/MRI is utilized for the diagnosis of gastrointestinal diseases, with each imaging modality being sensitive to different features. Different diseases necessitate attention to various imaging modalities, with some having established gold standards for diagnosis, such as WLE for Lipoma. However, many diseases lack a definitive gold standard  (Richardson et al. 1995; Chandler et al. 2019; Jacobson et al. 2023). The increasing medical focus on establishing diagnostic gold standards aims to standardize procedures and decrease required imaging modalities. The study seeks to provide evidence for definitive gold standards and serves as a reference for diseases without them. It is the pioneering research in computer-aided diagnosis to investigate the interpretability of detailed features and gold standard modalities, with the goal of enhancing AI’s role in setting diagnostic standards.
Fig. 3 illustrates the average modal attention of the proposed method for different disease categories. Firstly, the model exhibits notable specificity in modal attention for different disease categories. It is important to note that no constraints were imposed during model training regarding which modalities should be attended to for the recognition of different diseases. We attribute this specificity to the proposed modality weight adaptation module working in conjunction with multi-disease classification. Furthermore, we analyzed the alignment between the model’s specificity in modal attention for different disease categories and clinical priors. According to a study on gastrointestinal diseases by Jacobson et al. (Jacobson et al. 2023), using WLE alone to diagnose lipomas achieves a specificity of 99% in clinical practice. Similar findings hold true for NET. Conversely, GIST and EP rely more heavily on EUS. This alignment between the MDA weight and clinical expertise is highly consistent (Fig. 3), providing interpretability of modal attention for each disease category.
Table 6: Variations in adaptive weights for different modalities in the absence of certain modalities and learning with intrinsic noise.
Disease Modal Baseline w/o EUS w/o WLE w/o Report
GIST EUS 0.08$\pm$0.05 0.11$\pm$0.06 0.06$\pm$0.06 0.27$\pm$0.12
WLE 0.21$\pm$0.11 0.39$\pm$0.14 0.05$\pm$0.06 0.40$\pm$0.15
Report 0.72$\pm$0.13 0.49$\pm$0.13 0.90$\pm$0.10 0.33$\pm$0.13
GLM EUS 0.17$\pm$0.10 0.08$\pm$0.05 0.31$\pm$0.17 0.37$\pm$0.16
WLE 0.26$\pm$0.12 0.40$\pm$0.13 0.12$\pm$0.05 0.46$\pm$0.15
Report 0.56$\pm$0.14 0.52$\pm$0.11 0.56$\pm$0.14 0.17$\pm$0.09
NET EUS 0.33$\pm$0.10 0.11$\pm$0.04 0.26$\pm$0.10 0.71$\pm$0.13
WLE 0.10$\pm$0.07 0.35$\pm$0.11 0.10$\pm$0.05 0.16$\pm$0.11
Report 0.57$\pm$0.10 0.54$\pm$0.10 0.64$\pm$0.10 0.13$\pm$0.07
EP EUS 0.16$\pm$0.08 0.07$\pm$0.07 0.29$\pm$0.12 0.42$\pm$0.15
WLE 0.38$\pm$0.17 0.34$\pm$0.17 0.17$\pm$0.07 0.49$\pm$0.16
Report 0.46$\pm$0.16 0.59$\pm$0.16 0.54$\pm$0.11 0.09$\pm$0.06
Lipoma EUS 0.10$\pm$0.06 0.02$\pm$0.01 0.02$\pm$0.02 0.61$\pm$0.14
WLE 0.04$\pm$0.03 0.09$\pm$0.07 0.01$\pm$0.00 0.15$\pm$0.11
Report 0.86$\pm$0.07 0.89$\pm$0.07 0.97$\pm$0.02 0.24$\pm$0.11
GS EUS 0.35$\pm$0.13 0.13$\pm$0.08 0.33$\pm$0.12 0.77$\pm$0.09
WLE 0.27$\pm$0.13 0.33$\pm$0.13 0.27$\pm$0.06 0.12$\pm$0.08
Report 0.39$\pm$0.13 0.54$\pm$0.13 0.40$\pm$0.14 0.11$\pm$0.06
PC EUS 0.18$\pm$0.08 0.07$\pm$0.03 0.12$\pm$0.06 0.62$\pm$0.12
WLE 0.15$\pm$0.07 0.33$\pm$0.13 0.06$\pm$0.05 0.25$\pm$0.09
Report 0.67$\pm$0.12 0.59$\pm$0.12 0.82$\pm$0.05 0.13$\pm$0.12
IFP EUS 0.14$\pm$0.08 0.04$\pm$0.04 0.09$\pm$0.07 0.57$\pm$0.16
WLE 0.13$\pm$0.10 0.19$\pm$0.10 0.02$\pm$0.02 0.31$\pm$0.15
Report 0.72$\pm$0.13 0.77$\pm$0.10 0.89$\pm$0.08 0.12$\pm$0.06
Micro-visualization of model adaptation to missing modalities Table 6 demonstrates the variations in attention weights for three modalities before and after incorporating missing modalities and intrinsic noise in the network. The findings indicate that the proposed model is highly sensitive to these changes, as evidenced by the rapid and substantial shifts in attention weights across modalities when dealing with missing modalities. When EUS is missing, a significant decrease in EUS weights is first observed (0.34 to 0.05, 0.53 to 0.11), which is also observed when WLE is missing. When WLE is missing, attention weights almost entirely shift towards the report modality, which is expected since the surface information of the tumor, reflected in WLE, cannot be found in EUS. When WLE is absent, the model cannot learn relevant knowledge solely from the EUS modality. However, the report modality contains descriptions of WLE and can serve as a substitute for the missing image modality. Hence, the weights assigned to the report modality increase. Similarly, in the absence of EUS, the emphasis shifts more towards the reporting modality rather than WLE modality. In the scenario where the report modality is missing, the redistribution of the reduced weight is approximately evenly allocated to both WLE and EUS to compensate for the absence. Due to the lack of information provided by the missing modality, experiments indicate that when a missing modality is present (i.e., when the modality does not provide information), the MDA module can effectively reduce the impact of the missing modality on the final results through weight redistribution.
Conclusion
This paper introduces MDA, which employs a continuous attention mechanism to capture the linear correlations between multiple modalities. MDA exhibits the capability to effectively handle multi-modal information, even when dealing with missing modalities and intrinsic noise, eliminating the need for separate solutions for each scenario. To the best of our knowledge, This work is the first to explore intrinsic noise in diagnostic texts and MDA can withstand this noise interference. We have achieved SOTA results across various tasks and datasets, and have provided novel insights into the interpretability of modalities.
The interpretability of MDA in medical diagnosis is comprehensively analyzed. At the macroscopic level, we investigate the attention specificity of MDA towards different disease categories, demonstrating its alignment with clinical experience. At the microscopic level, we examine the significant changes in attention exhibited by MDA when handling missing modalities. The experimental results provide evidence of the efficacy of MDA in improving multi-modal fusion and its robustness in the presence of missing modalities or intrinsic noise. Additionally, we have made available a multi-modal dataset of gastrointestinal stromal tumors for future research purposes.